# Comparison of Lumerical FDTD and Tidy3D for three-dimensional FDTD simulations of passive silicon photonic components

### v2

@@ 2 Simulation configurations
Simulation of the same set of devices are conducted using Lumerical FDTD and Tidy3D under matched settings for comparison. The Lumerical FDTD solver version is 2025 R1 (8.33.3999), running locally using CPU on a workstation with AMD 3960X 24-core processor, 64 GB RAM, and Windows 10 OS. Tidy3D simulations are executed on Flexcompute’s GPU-accelerated cloud cluster using version 2.7.7 [13]. The computation resources are dynamically allocated based on simulation size and hardware availability at the time. After inquiry with Flexcompute, we were informed that one representative simulation - mode converter at the resolution of 25 cells per wavelength - was allocated to eight NVIDIA A100 GPUs. It needs to be noted that direct wall-clock comparisons reflect the best-effort use of each platform rather than identical compute resources. For both solvers, the geometry and simulation domain setup, boundary settings, and mesh configuration are all handled through their Python application programming interfaces (APIs). The Python codes are included in [14]. The auto-shutoff threshold is set to 110-5 for early termination. The simulation domain was defined to enclose the component with a distance of 2 m from its top and bottom edges and 1 m from edges on the device plane. Each device is extended with 10 m long straight waveguides at every port before importing into the solvers. This preprocessing step is to ensure extrusion of input and output waveguides through the boundaries in the propagation direction and therefore a reliable result. Perfectly matched layer (PML) boundary condition is applied to all six sides of the 3D domain. No symmetry planes are used to maintain generality. If a diverging simulation is encountered, we change the boundary condition accordingly. For instance, we switch to “absorber” in Tidy3D, which applies adiabatic absorbers containing multiple layers with gradually increasing conductivity to the boundaries. It is suitable for dispersive materials intersecting with simulation boundaries [15].
@@ 3.1 Directional coupler
The component is simulated at a series of resolutions ranging from 6 to 25 cells per wavelength. We notice that Tidy3D needs the boundary conditions to be set as “absorber” on all boundaries to avoid diverging simulation when the spatial resolution is 20 and 25. Table 2 lists the runtime in both solvers and the peak memory usage in Lumerical FDTD, with a source bandwidth of 20 nm centered at 1550 nm. The peak memory usage in Lumerical FDTD scales rapidly with increasing resolution due to finer discretization. With the same size of the simulation domain (4274 m3), the runtime of Tidy3D is much shorter than that of Lumerical FDTD. For example, Tidy3D completes the simulation with a resolution of 25 in roughly 10 minutes, while Lumerical FDTD needs almost 81 minutes. For this resolution, the total number of grid points and the elapsed simulation time and iterations are listed in Table 3. Tidy3D shows slightly finer discretization and runs for almost three times longer simulation time.
Spatial resolution | Runtime, | Runtime, | Memory usage, |
(cells per wavelength) | Tidy3D (s) | Lumerical FDTD (s) | Lumerical FDTD (GB) |
6 | 17 | 148 | 0.35 |
10 | 35 | 336 | 0.92 |
15 | 66 | 1293 | 2.42 |
20 | 136 | 2606 | 5.01 |
25 | 617 | 4856 | 9.07 |
@@ 3.2 Waveguide crossing
We examine simulations of this waveguide crossing in a series of resolutions from 6 to 25 cells per wavelength with a source bandwidth of 20 nm. The runtime and peak memory usage are listed in Table 4. The memory usage is lower than that for directional coupler simulations due to a smaller simulation domain. As expected, Tidy3D is much faster than Lumerical FDTD run on our workstation. With a simulation domain of 10114 m3 and a resolution of 25 cells per wavelength, Tidy3D completes the simulation in 1 minute and Lumerical completes it in 21 minutes. With this resolution, the total number of grid points and the elapsed simulation time and iterations are listed in Table 5. Both solvers show similar discretization, yet Tidy3D elapsed almost two times longer simulation time before the simulation reaches the auto-shutoff value.
Spatial resolution | Runtime, | Runtime, | Memory usage, |
(cells per wavelength) | Tidy3D (s) | Lumerical FDTD (s) | Lumerical FDTD (GB) |
6 | 11 | 31 | 0.23 |
10 | 18 | 85 | 0.51 |
15 | 28 | 234 | 1.26 |
20 | 84 | 591 | 2.57 |
25 | 61 | 1268 | 4.55 |
@@
The MMI is examined in both solvers with the same set of resolutions as in the previous sections with a source bandwidth of 20 nm centered at 1550 nm. Table 6 lists the runtime of the both solvers and the peak memory usage in Lumerical FDTD. As expected, Tidy3D is faster than Lumerical FDTD, completing the simulation with 25 cells per wavelength in 2 minutes 30 seconds while Lumerical needs 80 minutes. The simulation domain size is 347.54 m3. Table 7 lists the total number of grid points and the elapsed simulation time and iterations with the resolution of 25. Both solvers show similar discretization while Tidy3D elapsed a 30 longer simulation time.
Spatial resolution | Runtime, | Runtime, | Memory usage, |
(cells per wavelength) | Tidy3D (s) | Lumerical FDTD (s) | Lumerical FDTD (GB) |
6 | 11 | 98 | 0.32 |
10 | 34 | 259 | 0.82 |
15 | 45 | 905 | 2.15 |
20 | 57 | 2182 | 4.48 |
25 | 150 | 4813 | 8.10 |
@@ 3.4 Mode converter
The simulation is run at a series of spatial resolutions ranging from 6 to 25 cells per wavelength with a source bandwidth of 20 nm, as in the previous sections. The runtime in both solvers and the peak memory usage is listed in Table 8. Similar to the simulation of previous components, Tidy3D completes the job in much shorter time. For instance, Tidy3D finishes the simulation with the highest resolution (25 cells per wavelength) in 33 minutes while Lumerical FDTD needs 25 hours with a peak memory usage of 23.63 GB. We notice that the runtime of Tidy3D does not increase monotonically with spatial resolution, as the computational speed depends on the server resource usage at the time. The memory usage is more demanding than that in simulations of previous components due to a larger simulation domain (7210.54 m3). With the resolution of 25, the total number of grid points and the elapsed simulation time and iterations are listed in Table 9. Tidy3D shows finer discretization and elapses slightly longer simulation time.
Spatial resolution | Runtime, | Runtime, | Memory usage, |
(cells per wavelength) | Tidy3D (s) | Lumerical FDTD (s) | Lumerical FDTD (GB) |
6 | 63 | 760 | 0.71 |
10 | 517 | 3599 | 2.15 |
15 | 415 | 14417 | 6.00 |
20 | 124 | 40710 | 12.90 |
25 | 1997 | 91484 | 23.63 |
@@ 3.5 Polarization splitter rotator
The PSR is simulated at a range of resolutions from 6 to 25 cells per wavelength. Table 10 lists the runtime of the two solvers and the peak memory usage of Lumerical FDTD. Again, Tidy3D completes the simulations faster than Lumerical FDTD. In this case, the simulation domain size is 73.28.94 m3. For the highest resolution of 25 cells per wavelength, Tidy3D takes 21 minutes, while Lumerical FDTD needs 6 hours with 24 GB of memory usage. The total number of grid points and the elapsed simulation time and iterations are listed in Table 11. Tidy3D shows finer discretization and elapses almost two times longer simulation time.
Spatial resolution | Runtime, | Runtime, | Memory usage, |
(cells per wavelength) | Tidy3D (s) | Lumerical FDTD (s) | Lumerical FDTD (GB) |
6 | 22 | 303 | 0.66 |
10 | 66 | 1073 | 2.14 |
15 | 285 | 3478 | 6.05 |
20 | 604 | 9665 | 13.06 |
25 | 1262 | 21896 | 24.01 |
@@ 3.6 Ring resonator
We simulate the ring resonator for spatial resolutions varied from 6 to 25 cells per wavelength. The simulation domain size for the ring resonator is 3224.054 m3. The runtime in the two solvers and the peak memory usage in Lumerical FDTD are listed in Table 12. As for previous components, Tidy3D is faster than Lumerical FDTD. For instance, Tidy3D completes the highest resolution simulation in 42 minutes, while Lumeical FDTD requires 28 hours on our workstation. Table 13 lists the number of grids and the elapsed simulation time and iterations at the highest resolution of 25 cells per wavelength. Tidy3D applies slightly finer discretization along all three dimensions.
Spatial resolution | Runtime, | Runtime, | Memory usage, |
(cells per wavelength) | Tidy3D (s) | Lumerical FDTD (s) | Lumerical FDTD (GB) |
6 | 135 | 1074 | 3.14 |
10 | 236 | 4527 | 8.36 |
15 | 971 | 16524 | 18.50 |
20 | 2414 | 44034 | 32.62 |
25 | 2503 | 100421 | 49.42 |
@@ 4 Discussion
In the previous section, we present the simulation results of a selection of passive components on SOI platform sourced from GDSFactory generic PDK, including directional coupler, waveguide crossing, 22 MMI, mode converter, a polarization splitter rotator (PSR), and a ring resonator. Through the simulation of directional coupler and waveguide crossing, we evaluate the propagation and coupling of fundamental modes. The multimode behavior is examined by the simulation of MMI. Moreover, the simulation of mode converter and PRS demonstrates higher-order mode coupling and polarization transitioning. Lastly, we compare the resonance behavior of ring resonator between the two solvers. The minimum feature size across these components is 200 nm, occurring in the MMI between the tapers and in the ring resonator between the ring and the straight waveguide. In general, Tidy3D shows finer discretization given the same spatial resolution setting and elapses longer simulation time before reaching the auto-shutoff threshold.
@@
Component | Metrics | Lumerical FDTD | Tidy3D | ||
diectional coupler |
15 | 15 | ||
Simulation time | 22 min | 66 sec | |||
Bandwidth-induced variation | 3 | 0.2 | |||
Inter-solver discrepancy | 3.7 | ||||
waveguide crossing |
15 | 10 | ||
Simulation time | 4 min | 18 sec | |||
Bandwidth-induced variation | 0.1 | 0.03 | |||
Inter-solver discrepancy | 0.2 | ||||
22 MMI |
15 | 15 | ||
Simulation time | 15 min | 45 sec | |||
Bandwidth-induced variation | 0.05 | 0.05 | |||
Inter-solver discrepancy | 0.5 | ||||
mode converter |
10 | 15 | ||
Simulation time | 60 min | 7 min | |||
Bandwidth-induced variation | 9 | 3 | |||
Inter-solver discrepancy | 16 | ||||
PRS |
15 | 20 | ||
Simulation time | 58 min | 10 min | |||
Bandwidth-induced variation | 6 | 0.1 | |||
Inter-solver discrepancy | 2 | ||||
ring resonator |
20 | 20 | ||
Simulation time | 734 min | 40 min | |||
Bandwidth-induced variation | 47 | 0.5 | |||
Inter-solver discrepancy | 2 |
@@
Both Lumerical FDTD and Tidy3D are proven, reliable solvers for FDTD simulations. Lumerical FDTD is a mature solver offering extensive features and configurable settings. It can run efficiently on local hardware. The software includes a comprehensive example library and a scripting language. It also offers Python plugins (e.g. lumapi, lumopt). It can also be used in combination with other multiphysics solvers and circuit simulator offered by Lumerical. Tidy3D uses cloud-native GPU acceleration for rapid simulations, features a clear Python API and extensive examples, and is under active development to enhance its multiphysics and circuit simulation capabilities.
@@ 5 Conclusions
Acknowledgment The authors thank Wesley Sacher for insightful discussions about the selection of devices and materials in this work. The authors also thank Flexcompute for providing information on the latest update and computation hardware of Tidy3D.

### v3

@@ 2 Simulation configurations
Simulation of the same set of devices are conducted using Lumerical FDTD and Tidy3D under matched settings for comparison. The simulation results from Lumerical FDTD are obtained from a solver version of 2025 R1 (8.33.3999), running locally using CPU on a workstation with AMD 3960X 24-core processor, 64 GB RAM, and Windows 10 OS. Meanwhile, to demonstrate the performance on hardware more advanced than our local workstation, we also include the simulation runtime from Lumerical FDTD 2025 R2.1 that supports Ansys Cloud Burst Compute, which uses Amazon EC2 G6e instances powered by up to eight NVIDIA L40S GPUs. Tidy3D simulations are executed on Flexcompute’s GPU-accelerated cloud cluster using version 2.7.7 [13]. The computation resources are dynamically allocated based on simulation size and hardware availability at the time. After inquiry with Flexcompute, we were informed that one representative simulation - mode converter at the resolution of 25 cells per wavelength - was allocated to eight NVIDIA A100 GPUs. It needs to be noted that direct wall-clock comparisons reflect the best-effort use of each platform rather than identical compute resources. For both solvers, the geometry and simulation domain setup, boundary settings, and mesh configuration are all handled through their Python application programming interfaces (APIs). The Python codes are included in Code 1 [14]. The auto-shutoff threshold is set to 110-5 for early termination. The simulation domain was defined to enclose the component with a distance of 2 m from its top and bottom edges and 1 m from edges on the device plane. Each device is extended with 10 m long straight waveguides at every port before importing into the solvers. This preprocessing step is to ensure extrusion of input and output waveguides through the boundaries in the propagation direction and therefore a reliable result. Perfectly matched layer (PML) boundary condition is applied to all six sides of the 3D domain. No symmetry planes are used to maintain generality. If a diverging simulation is encountered, we change the boundary condition accordingly. For instance, we switch to “absorber” in Tidy3D, which applies adiabatic absorbers containing multiple layers with gradually increasing conductivity to the boundaries. It is suitable for dispersive materials intersecting with simulation boundaries [15].
@@ 3.1 Directional coupler
The component is simulated at a series of resolutions ranging from 6 to 25 cells per wavelength. We notice that Tidy3D needs the boundary conditions to be set as “absorber” on all boundaries to avoid diverging simulation when the spatial resolution is 20 and 25. Table 2 lists the runtime in both solvers and the peak memory usage in Lumerical FDTD, with a source bandwidth of 20 nm centered at 1550 nm. The peak memory usage in Lumerical FDTD scales rapidly with increasing resolution due to finer discretization. With the same size of the simulation domain (4274 m3), the runtime using advanced GPUs is much shorter than using our local workstation. The two solvers show similar runtime when similar hardware is provided. For this resolution, the total number of grid points and the elapsed simulation time and iterations are listed in Table 3. Tidy3D shows slightly finer discretization and runs for almost three times longer simulation time.
| Resolution | Runtime, | Local runtime, | Local memory usage, | Cloud runtime, |
| (cells/) | Tidy3D (s) | Lumerical FDTD (s) | Lumerical FDTD (GB) | Lumerical FDTD (s) |
| 6 | 18 | 148 | 0.35 | 13 |
| 10 | 36 | 356 | 0.92 | 19 |
| 15 | 63 | 1293 | 2.42 | 42 |
| 20 | 150 | 2606 | 5.01 | 77 |
| 25 | 626 | 4856 | 9.07 | 158 |
@@ 3.2 Waveguide crossing
We examine simulations of this waveguide crossing in a series of resolutions from 6 to 25 cells per wavelength with a source bandwidth of 20 nm. The runtime and peak memory usage are listed in Table 4. The memory usage is lower than that for directional coupler simulations due to a smaller simulation domain. As expected, Tidy3D is much faster than Lumerical FDTD run on our workstation. With a simulation domain of 10114 m3 and a resolution of 25 cells per wavelength, Tidy3D completes the simulation in 1 minute, while Lumerical completes it in 40 seconds on the cloud and 21 minutes locally. With this resolution, the total number of grid points and the elapsed simulation time and iterations are listed in Table 5. Both solvers show similar discretization, yet Tidy3D elapsed almost two times longer simulation time before the simulation reaches the auto-shutoff value.
| Spatial resolution | Runtime, | Local runtime, | Local memory usage, | Cloud runtime |
| (cells/) | Tidy3D (s) | Lumerical FDTD (s) | Lumerical FDTD (GB) | Lumerical FDTD (s) |
| 6 | 10 | 31 | 0.23 | 9 |
| 10 | 19 | 85 | 0.51 | 10 |
| 15 | 29 | 234 | 1.26 | 15 |
| 20 | 104 | 591 | 2.57 | 23 |
| 25 | 78 | 1268 | 4.55 | 39 |
@@
The MMI is examined in both solvers with the same set of resolutions as in the previous sections with a source bandwidth of 20 nm centered at 1550 nm. Table 6 lists the runtime of the both solvers and the peak memory usage in Lumerical FDTD. The simulation domain size is 347.54 m3. Table 7 lists the total number of grid points and the elapsed simulation time and iterations with the resolution of 25. Both solvers show similar discretization while Tidy3D elapsed a 30 longer simulation time.
| Resolution | Runtime, | Local runtime, | Local memory usage, | Cloud runtime, |
| (cells/) | Tidy3D (s) | Lumerical FDTD (s) | Lumerical FDTD (GB) | Lumerical FDTD (s) |
| 6 | 11 | 98 | 0.32 | 12 |
| 10 | 34 | 259 | 0.82 | 15 |
| 15 | 45 | 905 | 2.15 | 34 |
| 20 | 63 | 2182 | 4.48 | 72 |
| 25 | 163 | 4813 | 8.10 | 119 |
@@ 3.4 Mode converter
The simulation is run at a series of spatial resolutions ranging from 6 to 25 cells per wavelength with a source bandwidth of 20 nm, as in the previous sections. The runtime in both solvers and the peak memory usage is listed in Table 8. We notice that the runtime of Tidy3D does not increase monotonically with spatial resolution, as the computational speed depends on the server resource usage at the time. The memory usage is more demanding than that in simulations of previous components due to a larger simulation domain (7210.54 m3). With the resolution of 25, the total number of grid points and the elapsed simulation time and iterations are listed in Table 9. Tidy3D shows finer discretization and elapses slightly longer simulation time.
| Resolution | Runtime, | Local runtime, | Local memory usage, | Cloud runtime, |
| (cells/) | Tidy3D (s) | Lumerical FDTD (s) | Lumerical FDTD (GB) | Lumerical FDTD (s) |
| 6 | 63 | 750 | 0.71 | 30 |
| 10 | 518 | 3599 | 2.15 | 108 |
| 15 | 413 | 14417 | 6.00 | 393 |
| 20 | 133 | 40710 | 12.90 | 384 |
| 25 | 2011 | 91484 | 23.63 | 810 |
@@ 3.5 Polarization splitter rotator
The PSR is simulated at a range of resolutions from 6 to 25 cells per wavelength. Table 10 lists the runtime of the two solvers and the peak memory usage of Lumerical FDTD. Again, Tidy3D completes the simulations faster than Lumerical FDTD. In this case, the simulation domain size is 73.28.94 m3. The total number of grid points and the elapsed simulation time and iterations are listed in Table 11. Tidy3D shows finer discretization and elapses almost two times longer simulation time.
| Resolution | Runtime, | Local runtime, | Local memory usage, | Cloud runtime, |
| (cells/) | Tidy3D (s) | Lumerical FDTD (s) | Lumerical FDTD (GB) | Lumerical FDTD (s) |
| 6 | 22 | 303 | 0.66 | 25 |
| 10 | 65 | 1073 | 2.14 | 58 |
| 15 | 81 | 3478 | 6.05 | 150 |
| 20 | 613 | 9665 | 13.06 | 277 |
| 25 | 1271 | 21896 | 24.01 | 481 |
@@ 3.6 Ring resonator
We simulate the ring resonator for spatial resolutions varied from 6 to 25 cells per wavelength. The simulation domain size for the ring resonator is 3224.054 m3. The runtime in the two solvers and the peak memory usage in Lumerical FDTD are listed in Table 12. As for previous components, Tidy3D is faster than Lumerical FDTD. Table 13 lists the number of grids and the elapsed simulation time and iterations at the highest resolution of 25 cells per wavelength. Tidy3D applies slightly finer discretization along all three dimensions.
| Resolution | Runtime, | Local runtime, | Local memory usage, | Cloud runtime, |
| (cells/) | Tidy3D (s) | Lumerical FDTD (s) | Lumerical FDTD (GB) | Lumerical FDTD (s) |
| 6 | 233 | 1168 | 3.14 | 52 |
| 10 | 307 | 4685 | 8.36 | 157 |
| 15 | 1104 | 16815 | 18.50 | 476 |
| 20 | 2594 | 44530 | 32.62 | 732 |
| 25 | 2819 | 101191 | 49.42 | 1433 |
@@ 4 Discussion
In the previous section, we present the simulation results of a selection of passive components on SOI platform sourced from GDSFactory generic PDK, including directional coupler, waveguide crossing, 22 MMI, mode converter, a polarization splitter rotator (PSR), and a ring resonator. Through the simulation of directional coupler and waveguide crossing, we evaluate the propagation and coupling of fundamental modes. The multimode behavior is examined by the simulation of MMI. Moreover, the simulation of mode converter and PRS demonstrates higher-order mode coupling and polarization transitioning. Lastly, we compare the resonance behavior of ring resonator between the two solvers. The minimum feature size across these components is 200 nm, occurring in the MMI between the tapers and in the ring resonator between the ring and the straight waveguide. In general, Tidy3D shows finer discretization given the same spatial resolution setting and elapses longer simulation time before reaching the auto-shutoff threshold. The two solvers give similar runtime when run on advanced NVIDIA GPUs. It needs to be noted that the latest version of Lumerical FDTD (R2.1) does not support frequency dependent mode profiles when solving on GPUs. This could potentially impact broadband simulations.
@@
| Component | Metrics | Lumerical FDTD | Tidy3D |
| diectional coupler | Required resolution (cells per wavelength) | 15 | 15 |
| Bandwidth-induced variation | 3 | 0.2 | |
| Inter-solver discrepancy | 3.7 | ||
| waveguide crossing | Required resolution (cells per wavelength) | 15 | 10 |
| Bandwidth-induced variation | 0.1 | 0.03 | |
| Inter-solver discrepancy | 0.2 | ||
| 22 MMI | Required resolution (cells per wavelength) | 15 | 15 |
| Bandwidth-induced variation | 0.05 | 0.05 | |
| Inter-solver discrepancy | 0.5 | ||
| mode converter | Required resolution (cells per wavelength) | 10 | 15 |
| Bandwidth-induced variation | 9 | 3 | |
| Inter-solver discrepancy | 16 | ||
| PRS | Required resolution (cells per wavelength) | 15 | 20 |
| Bandwidth-induced variation | 6 | 0.1 | |
| Inter-solver discrepancy | 2 | ||
| ring resonator | Required resolution (cells per wavelength) | 20 | 20 |
| Bandwidth-induced variation | 47 | 0.5 | |
| Inter-solver discrepancy | 2 | ||
@@
Both Lumerical FDTD and Tidy3D are proven, reliable solvers for FDTD simulations. Lumerical FDTD is a mature solver offering extensive features and configurable settings. It can run efficiently on local and cloud-based hardware. The software includes a comprehensive example library and a scripting language. It also offers Python plugins (e.g. lumapi, lumopt). It can also be used in combination with other multiphysics solvers and circuit simulator offered by Lumerical. Tidy3D uses cloud-native GPU acceleration for rapid simulations, features a clear Python API and extensive examples, and is under active development to enhance its multiphysics and circuit simulation capabilities.
@@ 5 Conclusions
Acknowledgment The authors thank Wesley Sacher for insightful discussions about the selection of devices and materials in this work. The authors also thank Flexcompute for providing information on the latest update and computation hardware of Tidy3D. Additionally, the authors thank Ansys for providing access to Ansys Cloud Burst Compute and discussion about definition of runtime.